OpenStudy (anonymous):

Carlos solved the system of equations below. What mistake did he make in his work? 2x + y = 5 x − 2y = 10 y = 5 − 2x x − 2(5 − 2x) = 10 x + 10 + 4x = 10 5x + 10 = 10 5x = 0 x = 0 2(0) + y = 5 y = 5

OpenStudy (anonymous):

wrong expansion x − 2(5 − 2x) = 10 x - 10 + 4x = 10 not x + 10 + 4x = 10